the line join of the points(2,3) and (-1,5) subtends a right angle at P. If the locus of P is a circle then it`s radius is

Grade:11

1 Answers

Arun
6 years ago
Dear Kamal
Let the P = (x, y)
Now
Slope of PA = (y-3)/(x-2)
And slope of PB = (y-5)/(x+1)
It's multiply should be -1.
[(y-3)/(x-2)] *[(y-5)/(x+1)] = -1
Now expand it ,
You will get a equation like a circle and then you can also find the radius.
 
Regards
Arun(askIITians forum expert)